OpenStudy (anonymous):

Anais bought 3 1/2 yards of ribbon. She had 2 feet 10 inches of ribbon left after trimming some curtains. How many inches of ribbon did Anais use to trim the curtains? Anais used _____ inches of ribbon. The solution is .

OpenStudy (anonymous):

convert all in inches... 1 yd = 3 feet 1 foot = 12 inches

OpenStudy (anonymous):

\[3.5 \cancel{yards} \times \frac{3 \cancel{feet}}{1 \cancel{yard}} \times \frac{12~inches}{1 \cancel{foot}}=126~inches\]

OpenStudy (anonymous):

2 feet = 24 inches 24 + 10 = 34 inches remains after trimming some curtains

OpenStudy (anonymous):

Anais used... 126 inches - 34 inches = 92 inches // Answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hope you understand the solution... \(\ddot\smile\)

OpenStudy (anonymous):

i do